# Investigating Context-Faithfulness in Large Language Models: The Roles of Memory Strength and Evidence Style

### v1

@@ 3.4.2 Evidence Generation
To ensure the reliability of the evidence, the evidence must mutually entail with the CMA. This entailment is verified using an NLI model  [^4] https://huggingface.co/microsoft/deberta-v2-xxlarge-mnli. Direct evidence is intuitive, simple, and coherent, making it the straightforward type of evidence for the LLM to process. If the LLM is receptive to external evidence, it should be able to easily adopt direct evidence.
@@
To ensure the reliability of indirect evidence, the indirect evidence should entail the CMA, and the additional information introduced by the evidence should not entail the MA. Otherwise, the indirect evidence can support both the MA and CMA. The NLI model4 is used to verify that indirect evidence entailed with CMA and neutral or contradictory with MA.
@@ 4 Experiments
In this study, we aim to investigate two key research questions. 1) Whether memory strength has an impact on the context faithfulness of LLMs. 2) If the style of evidence affects the context faithfulness of LLMs. These research questions are explored in Section 4.2 and 4.3, respectively. We also provide additional studies in Appendix A, which includes a study about the impact of option order and a case study.
@@ 4.1 Experiment Setup
LLM Models. Our experiments were conducted using four well-known language models: ChatGPT OpenAI (2023a), GPT-4 OpenAI (2023b), LLaMA2.7B, and LLaMA2.70B Touvron et al. (2023). These models represent a diverse range of architectures and capabilities. ChatGPT and GPT-4 are cutting-edge models developed by OpenAI, with GPT-4 being the most advanced in terms of reasoning, accuracy, and contextual understanding. LLaMA2, with its 7 billion and 70 billion parameter variants, is a strong open-source alternative that has demonstrated competitive performance across a wide variety of tasks. The inclusion of models with varying scales (from 7B to 70B) and training methodologies allows us to explore both closed-source systems (GPT-4 and ChatGPT) and open-source solutions (LLaMA2).
Evaluation Metrics. Following previous work Longpre et al. (2021); Xie et al. (2024); Chen et al. (2022), we transform the free-form QA to a multiple-choice QA format by providing a few options as possible answers [^5] Xie et al. (2024) shows that answer consistency between free-form and multi-choice are 94%, 96% and 92% for ChatGPT, GPT-4 and LLaMA2.7B, respectively.. This limits the generation space and helps determine the answer provided by LLMs with certainty. Specifically, for each question from both datasets, LLMs are instructed to select one answer from MA, CMA, and “Uncertain” (UCT). We report the ratio of MA, CMA, and UCT, and the formulas are
@@
where $f_{m},f_{c},f_{u}$ are the count of questions with MA, CMA, and UCT answers, respectively.
4.2 Impact of Memory Strength
4.2.1 Memory Strength on Different Datasets
@@
Figure 3: Relationship between Memory Strength and Answer Ratio across popQA and NQ datasets (with Direct Evidence). The figure presents the ratio and count of MA, CMA, and UCT across four memory strength groups: low(lo), mid-low(ml), mid-high(mh), and high(hi). The results show a clear positive correlation between memory strength and MA ratio($R_{m}$).
We first illustrate the distributions of memory strength across the popQA and NQ datasets for LLaMA2.7B, LLaMA2.70B, ChatGPT, and GPT-4, respectively (shown in Figure 2). From the analysis, two key insights can be drawn.
LLMs demonstrate stronger memory for the NQ dataset than the popQA dataset. For the NQ dataset, most questions fall within the bin of (0.25, 0]. Only a few questions fall within bins of weaker memory strength. In contrast, the popQA dataset has a greater number of questions in bins with weaker memory strength. This phenomenon is consistent across all four evaluated LLMs, indicating that the LLMs exhibit stronger memory strength for the NQ dataset compared to the PopQA dataset. A possible explanation is that the NQ dataset covers more commonly discussed subjects than those in the PopQA dataset. These subjects may have been encountered more frequently during the training of the LLMs, making it easier for the models to recall the information and resulting in stronger memory strength.
Memory Strength Increases with Model Scale. There is a clear trend that larger LLMs, such as GPT-4, demonstrate significantly higher memory strength compared to smaller models, such as LLaMA2.7B. It means larger models are more likely to provide consistent answers to paraphrased questions. This observation aligns with the common intuition that larger LLMs with more parameters have more memory than smaller LLMs.
4.2.2 Context-Faithfulness with Memory Strength
@@
To demonstrate the relationship between context-faithfulness and memory strength, we categorize the questions in each dataset into four groups according to memory strength for each LLM. The four groups are low, mid-low, mid-high, and high, corresponding to the memory strength intervals [-2, -1], (-1, -0.5], (-0.5, -0.25] and [-0.25, 0], respectively. We use the direct evidence. The results are shown in Figure 3666We put results for other evidence styles in Appendix, Figure 6, 7, 8. The conclusion is consistent.. Figure 3 (a)(b) shows the ratios of questions with MA, CMA, and UCT answers for the NQ and popQA datasets, respectively. Note that different LLMs may have different memory strengths to the same question. Therefore, the count of questions and the specific questions within the same group can vary across different LLMs. To illustrate this, we present the count of questions in each group (low, mid-low, mid-high, and high) in Figure 3 (c)(d) for popQA and NQ datasets, respectively. We can draw the following conclusions.
@@
There is a clear positive correlation between memory strength and MA ratio. From Figure 3 (a)(b), it is obvious that as memory strength increases, the ratio of MA (red) increases significantly, while the proportion of CMA (blue) decreases. This trend is consistent across both datasets, but it is more obvious in the NQ dataset, especially for larger models such as GPT-4 and ChatGPT.
In general, the LLaMA models are more context-faithful than the GPT models. Comparing the LLaMA models (7B and 70B) with the GPT models (ChatGPT and GPT-4) in Figure 3 (a)(b), the LLaMA models have a lower MA ratio and a higher CMA ratio. Even within groups of the same memory strength (low or high), the LLaMA models demonstrate more context-faithful performance.
4.3 Impact of Evidence Style
@@
LLaMA2.7B LLaMA2.70B ChatGPT GPT-4 LLaMA2.7B LLaMA2.70B ChatGPT GPT-4
##\## of Q (Initial) 1000 1000 1000 1000 1667 1667 1667 1667
##\## of Q with direct evidence 918 922 933 933 1042 1009 1079 1171
##\## of Q with indirect evidence 901 895 911 918 976 972 1025 1108
@@
Dataset Evidence Style S # LLaMA2.7B LLaMA2.70B ChatGPT GPT-4
$R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$
@@
NQ Direct 1 7.2 92.8 0.0 3.07 96.93 0.0 19.46 75.16 5.38 50.04 47.99 1.96
2 5.47 94.53 0.0 3.07 96.93 0.0 19.09 76.83 4.08 20.24 77.54 2.22
3 6.81 93.19 0.0 2.68 97.22 0.1 22.06 72.75 5.19 17.34 80.87 1.79
Direct + Paraphrase 2 4.13 95.87 0.0 1.49 98.41 0.1 15.29 81.28 3.43 18.96 79.67 1.37
3 3.26 96.74 0.0 1.19 98.61 0.2 9.55 86.75 3.71 11.27 87.28 1.45
@@
Direct 1 5.53 94.47 0.0 2.67 97.32 0.0 18.73 75.71 5.56 48.83 49.19 1.99
Indirect 2 3.28 95.29 1.43 1.65 98.25 0.1 13.66 84.1 2.24 44.59 53.7 1.71
3 4.82 94.06 1.13 1.85 97.84 0.31 13.27 84.19 2.54 39.89 58.57 1.53
Direct + Indirect 2 5.33 94.67 0.0 1.34 98.25 0.41 12.68 84.78 2.54 32.4 66.06 1.53
3 4.41 95.59 0.0 1.44 98.56 0.0 9.46 88.1 2.44 28.7 69.67 1.62
@@
popQA Direct 1 0.44 99.56 0.0 1.08 98.7 0.22 3.32 94.75 1.93 13.29 84.57 2.14
2 0.44 99.56 0.0 0.98 98.81 0.22 2.79 96.03 1.18 4.39 93.46 2.14
3 0.65 99.35 0.0 1.08 98.7 0.22 3.0 95.5 1.5 2.57 95.28 2.14
Direct + Paraphrase 2 0.22 99.78 0.0 1.19 98.7 0.11 2.36 97.0 0.64 3.0 95.71 1.29
3 0.11 99.89 0.0 0.43 99.35 0.22 1.39 98.28 0.32 1.29 98.5 0.21
@@
Direct 1 0.44 99.56 0.0 0.45 99.33 0.22 3.07 95.28 1.65 12.53 85.29 2.18
Indirect 2 0.0 100.0 0.0 0.11 99.89 0.0 3.18 96.38 0.44 13.51 85.73 0.76
3 0.0 100.0 0.0 0.0 100.0 0.0 1.76 97.91 0.33 9.26 90.2 0.55
Direct + Indirect 2 0.22 99.78 0.0 0.11 99.78 0.11 1.87 97.26 0.88 7.95 91.18 0.87
3 0.11 99.89 0.0 0.11 99.78 0.11 1.43 98.13 0.44 5.12 94.77 0.11
Table 2: Results of LLM Receptiveness to Different Evidence Styles Across NQ and popQA Datasets. The table presents the MA ratio ($R_{m}$), CMA ratio ($R_{c}$), and uncertain answer ratio ($R_{u}$) for various evidence styles across four models. All the ratios are in %.
Evidence Styles. We formulate four types of evidence styles: 1) Direct Evidence. This is the most straightforward form of evidence and serves as our baseline. To assess the impact of evidence length, we also create versions where the direct evidence is repeated twice and three times for comparison. 2) Direct Evidence Combined with Paraphrases of CMA (Direct + Paraphrase). To examine the effect of evidence phrasing and expression, we combine the direct evidence with one paraphrase of the CMA to form a two-sentence evidence and with two paraphrases to form a three-sentence evidence. 3) Indirect Evidence. We generate indirect evidence consisting of two sentences and three sentences, respectively [^7] We regulate the length of the generated evidence to control the influence of evidence length. The prompts used are detailed in Table 9 (index 7) in Appendix.. 4) Direct Evidence Combined with Indirect Evidence (Direct + Indirect). We combine the direct evidence with the first sentence of the two-sentence indirect evidence to form a two-sentence evidence and with both sentences to form a three-sentence evidence.
@@
In Group 2, $R_{m}$ of direct evidence with one sentence is slightly lower than that in Group 1. During the evidence generation, there are some questions for which ChatGPT can provide direct evidence but cannot produce indirect evidence. Removing these questions leads to a decrease in $R_{m}$ of direct evidence with one sentence, which implies that LLMs have a relatively high $R_{m}$ for the removed questions. But in general, the $R_{m}$ of direct evidence with one sentence in Group 1 is close to that in Group 2, so the results of Group 1 and Group 2 are still comparable.
Simple repetition of direct evidence is ineffective, except for GPT-4. Comparing direct evidence with one to three sentences, we observe similar $R_{m}$ and $R_{c}$ for most LLMs, except for GPT-4. On direct evidence with two and three sentences, $R_{m}$ of GPT-4 decreases significantly, and $R_{c}$ increases significantly. The results imply that simple repetition of direct evidence has no improvement in the LLMs’ receptiveness to CMAs for LLaMA2.7B, LLaMA2.70B, and ChatGPT.
Paraphrasing direct evidence is highly effective across all models and datasets. Comparing direct with two and three sentences and direct + paraphrase with two and three sentences, we observe $R_{m}$ of paraphrase evidence decrease significantly. For example, $R_{m}$ is reduced by more than half, comparing paraphrase with three sentences with direct with three sentences. The result implies that paraphrasing is an effective method to enhance the receptiveness of LLMs to external evidence.
Indirect evidence improves LLMs’ receptiveness to CMAs, but less effectively than paraphrasing. Comparing indirect evidence with two and three sentences with direct evidence with one sentence, $R_{m}$ decreased. But compared to the direct + paraphrase evidence with two or three sentences, the reduction is not significant. It implies that adding detailed information is less effective than paraphrasing direct evidence.
Combining direct evidence with indirect evidence generally enhances persuasiveness, except for LLaMA2.7B. Comparing direct + indirect evidence with indirect evidence, $R_{m}$ decreases except for LLaMA2.7B. For example, comparing direct + indirect with three sentences and indirect evidence with three sentences, $R_{m}$ has an obvious decrease. The result implies that adding direct evidence to indirect evidence is effective in improving LLMs’ receptiveness to CMAs.

### v2

@@ 3.4.2 Evidence Generation
To ensure the reliability of the evidence, the evidence must mutually entail with the CMA. This entailment is verified using an NLI model  [^4] https://huggingface.co/microsoft/deberta-v2-xxlarge-mnli. Direct evidence is intuitive, simple, and coherent, making it the straightforward type of evidence for the LLM to process. If the LLM is receptive to external evidence, it should be able to adopt direct evidence easily.
@@
To ensure the reliability of indirect evidence, the indirect evidence should entail the CMA and the additional information introduced by the evidence should not entail the MA. Otherwise, the indirect evidence can support both the MA and CMA. The NLI model4 is used to verify that indirect evidence entailed with CMA and neutral or contradictory with MA.
@@ 4 Experiments
In this study, we aim to investigate two key research questions. 1) Does memory strength have an impact on the context faithfulness of LLMs? 2) Does the style of evidence affect the context faithfulness of LLMs? These research questions are explored in Section 4.2 and 4.3, respectively. We also provide additional studies in Appendix B, which includes a study about the impact of option order and a case study.
@@ 4.1 Experiment Setup
LLM Models. Our experiments are conducted using six well-known language models: ChatGPT OpenAI (2023a), GPT-4 OpenAI (2023b), LLaMA2-7B, LLaMA2-70B Touvron et al. (2023), LLaMA3.2-3B Meta (2024), and Claude3.5 Anthropic (2024). These models represent a diverse range of architectures and capabilities. ChatGPT, GPT-4, and Claude3.5 are cutting-edge models developed by OpenAI and Anthropic. LLaMA2, with its 7 billion and 70 billion parameter variants, is a strong open-source alternative that has demonstrated competitive performance across a wide variety of tasks. LLaMA3.2-3B is the newest version of LLaMA. The inclusion of models with varying scales (from 3B to 70B) and training methodologies allows us to explore both closed-source systems (GPT-4, ChatGPT, and Claude3.5) and open-source solutions (LLaMA2-7B, LLaMA-70B and LLaMA3.2-3B).
Evaluation Metrics. Following previous work Longpre et al. (2021); Xie et al. (2024); Chen et al. (2022), we transform the short answer QA to a multiple-choice QA format by providing a few options as possible answers [^5] Xie et al. (2024) shows that answer consistency between short answer and multi-choice are 94%, 96% and 92% for ChatGPT, GPT-4 and LLaMA2-7B, respectively.. This limits the answer generation space and makes it easy to evaluate without manual checking. Specifically, for each question from both datasets, LLMs are instructed to select one answer from the MA, CMA, and “Uncertain” (UCT). We report the ratio of MA ($R_{m}$), CMA ($R_{c}$), and UCT ($R_{u}$) as calculated below:
@@
where $f_{m},f_{c}$, and $f_{u}$ are the count of questions with MA, CMA, and UCT answers, respectively.
4.2 Role of Memory Strength
@@
Figure 3: Relationship between Memory Strength and Different Answers’ Ratios across popQA and NQ Datasets (with Direct Evidence). The figure presents the ratio and count of MA, CMA, and UCT across four memory strength groups: low(lo), mid-low(ml), mid-high(mh), and high(hi). The results show a clear positive correlation between memory strength and MA ratio($R_{m}$).
4.2.1 Correlation between Context Faithfulness and Memory Strength
@@
To demonstrate the relationship between context faithfulness and memory strength, we categorize the questions in each dataset into four groups according to the memory strength of each LLM. The four groups are low, mid-low, mid-high, and high, corresponding to the memory strength intervals [-2, -1], (-1, -0.5], (-0.5, -0.25] and [-0.25, 0], respectively. We use the direct evidence in this experiment. The results are shown in Figure 3666We put results for other evidence styles in Appendix, Figures 8, 9, and 10. The conclusion is consistent.. Figure 3 (a)(b) shows the ratios of questions with MA, CMA, and UCT answers for the NQ and popQA datasets, respectively. Note that different LLMs may have different memory strengths to the same question. Therefore, both the specific questions and the count of questions in the same group can vary across different LLMs. To illustrate this, we present the count of questions in each group (low, mid-low, mid-high, and high) in Figure 3 (c)(d) for popQA and NQ datasets, respectively. We can draw the following conclusions.
@@
There is a clear positive correlation between memory strength and MA ratio for individual LLMs. From Figure 3 (a)(b), it is obvious that for all tested LLMs, the ratio of MA (red) increases when memory strength increases, while the ratio of CMA (blue) decreases. This trend is also consistent across both datasets. It is more obvious for GPT models. Among the tested LLMs, Claude3.5 tends to choose UCT options more often, especially on questions with high memory strength.
Memory Strength Increases with Model Scale. We can observe from Figure 3 (c)(d) that larger LLMs, such as GPT-4, have more data samples in the high memory strength group (hi), while smaller LLMs, such as LLaMA3.2-3B, have more samples in the low memory strength group (lo). This aligns with the common intuition that larger LLMs, with more parameters, have more knowledge than smaller LLMs. Further evidence and discussion can be found in the Appendix B.1.
4.2.2 LLMs Performance Analysis
Figure 4: MA Ratio ($R_{m}$) and UCT Ratio ($R_{u}$) VS Memory Strength across PopQA and NQ Datasets. The figure presents the relationship between average memory strength and the MA ratio as well as the UCT ratio for all six LLMs. It shows that lower average memory strength does not always mean better context faithfulness for different LLMs and newer versions of GPT and LLaMA models seem to ignore context faithfulness issues.
In the aforementioned conclusion, different LLMs have different knowledge. Thus, testing LLMs’ context faithfulness on the same dataset may be unfair. To illustrate this, we compute the average memory strength of LLMs on PopQA and NQ datasets, respectively, along with their MA and UCT ratios. The results are presented in Figure 4. We can draw the following conclusions.
First, for different LLMs, a lower average memory strength does not necessarily imply better context faithfulness. For example, Claude 3.5 has a high average memory strength, implying that Claude 3.5 is a knowledgeable LLM, but it has a low MA ratio ($R_{m}$) and a high UCT ratio ($R_{u}$). This indicates that Claude 3.5 tends to refuse to answer when facing knowledge conflicts. In contrast, LLaMA3.2-3B, despite having much less knowledge (low average memory strength), has the second-highest MA ratio ($R_{m}$). This means that LLaMA3.2-3B relies heavily on limited internal knowledge when facing knowledge conflicts, implying that it is not context-faithful.
Second, newer versions of GPT and LLaMA models appear to overlook context faithfulness issues during the training process. GPT-4, with slightly more knowledge than Chatgpt, shows a significantly higher MA ratio ($R_{m}$), and LLaMA3.2-3B, with the least knowledge, shows a higher MA ratio ($R_{m}$) than the two LLaMA2 models.
Finally, a new context faithfulness evaluation metric is needed. This metric should consider both the different answers’ ratios and memory strength when LLMs encounter knowledge conflicts. Simply using the MA ratio (which is used widely to measure context faithfulness currently) to evaluate context faithfulness across different LLMs may not be fair and sufficient.
4.3 Role of Evidence Style
@@
LLaMA2-7B LLaMA2-70B LLaMA3.2-3B ChatGPT GPT-4 Claude3.5 LLaMA2-7B LLaMA2-70B LLaMA3.2-3B ChatGPT GPT-4 Claude3.5
##\## of Q (Initial) 1000 1000 1000 1000 1000 1000 1667 1667 1667 1667 1667 1667
##\## of Q with direct evidence 918 922 938 933 933 931 1042 1009 1002 1079 1171 1173
##\## of Q with indirect evidence 901 895 917 911 918 913 976 972 941 1025 1108 1059
@@
Dataset Evidence Style S # LLaMA2-7B LLaMA2-70B LLaMA3.2-3B ChatGPT GPT-4 Claude3.5
$R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$
@@
popQA Direct Evidence 1 0.44 99.56 0.0 1.08 98.7 0.22 7.69 85.15 7.16 3.32 94.75 1.93 13.29 84.57 2.14 0.65 83.31 16.04
2 0.44 99.56 0.0 0.98 98.81 0.22 2.67 95.83 1.5 2.79 96.03 1.18 4.39 93.46 2.14 0.32 94.08 5.6
3 0.65 99.35 0.0 1.08 98.7 0.22 2.56 94.98 2.46 3.0 95.5 1.5 2.57 95.28 2.14 0.21 94.73 5.06
Direct+Paraphrase 2 0.22 99.78 0.0 1.19 98.7 0.11 3.1 95.62 1.28 2.36 97.0 0.64 3.0 95.71 1.29 0.21 95.69 4.09
3 0.11 99.89 0.0 0.43 99.35 0.22 1.07 97.86 1.07 1.39 98.28 0.32 1.29 98.5 0.21 0.11 98.06 1.83
@@
Direct Evidence 1 0.44 99.56 0.0 0.45 99.33 0.22 7.31 85.61 7.09 3.07 95.28 1.65 12.53 85.29 2.18 0.66 83.9 15.44
Indirect Evidence 2 0.0 100.0 0.0 0.11 99.89 0.0 4.36 89.64 6.0 3.18 96.38 0.44 13.51 85.73 0.76 0.88 82.37 16.76
3 0.0 100.0 0.0 0.0 100.0 0.0 3.27 92.58 4.14 1.76 97.91 0.33 9.26 90.2 0.55 0.88 87.95 11.17
Direct+Indirect 2 0.22 99.78 0.0 0.11 99.78 0.11 3.71 93.24 3.05 1.87 97.26 0.88 7.95 91.18 0.87 0.88 93.43 5.7
3 0.11 99.89 0.0 0.11 99.78 0.11 1.42 98.15 0.44 1.43 98.13 0.44 5.12 94.77 0.11 0.88 96.28 2.85
@@
NQ Direct Evidence 1 7.2 92.8 0.0 3.07 96.93 0.0 26.41 59.88 13.71 19.46 75.16 5.38 50.04 47.99 1.96 1.96 56.4 41.64
2 5.47 94.53 0.0 3.07 96.93 0.0 16.93 75.2 7.86 19.09 76.83 4.08 20.24 77.54 2.22 0.77 83.7 15.53
3 6.81 93.19 0.0 2.68 97.22 0.1 11.79 76.11 12.1 22.06 72.75 5.19 17.34 80.87 1.79 0.26 88.22 11.52
Direct+Paraphrase 2 4.13 95.87 0.0 1.49 98.41 0.1 13.21 79.33 7.46 15.29 81.28 3.43 18.96 79.67 1.37 0.34 88.65 11.01
3 3.26 96.74 0.0 1.19 98.61 0.2 9.38 83.27 7.36 9.55 86.75 3.71 11.27 87.28 1.45 0.26 93.09 6.65
@@
Direct Evidence 1 5.53 94.47 0.0 2.67 97.32 0.0 23.19 63.73 13.08 18.73 75.71 5.56 48.83 49.19 1.99 1.77 59.44 38.79
Indirect Evidence 2 3.28 95.29 1.43 1.65 98.25 0.1 13.32 77.77 8.92 13.66 84.1 2.24 44.59 53.7 1.71 1.98 67.88 30.14
3 4.82 94.06 1.13 1.85 97.84 0.31 11.53 80.86 7.61 13.27 84.19 2.54 39.89 58.57 1.53 3.65 71.32 25.03
Direct+Indirect 2 5.33 94.67 0.0 1.34 98.25 0.41 9.75 82.28 7.97 12.68 84.78 2.54 32.4 66.06 1.53 1.36 80.92 17.73
3 4.41 95.59 0.0 1.44 98.56 0.0 8.32 85.49 6.18 9.46 88.1 2.44 28.7 69.67 1.62 1.77 83.94 14.29
Table 3: Results of LLM Receptiveness to Different Evidence Styles Across NQ and popQA Datasets. The table presents the MA ratio ($R_{m}$), CMA ratio ($R_{c}$), and uncertain answer ratio ($R_{u}$) for various evidence styles across six models. All the ratios are in %. The best results are highlighted in bold, and the second-best results are underlined.
Evidence Styles. We formulate four types of evidence styles: 1) Direct Evidence. This is the most straightforward form of evidence and serves as our baseline. To assess the impact of evidence length, we also create versions where the direct evidence is repeated twice and three times for comparison. 2) Direct Evidence Combined with Paraphrases of CMA. To examine the effect of evidence phrasing and expression, we combine the direct evidence with one paraphrase of the CMA to form a two-sentence evidence and with two paraphrases to form a three-sentence evidence. 3) Indirect Evidence. We generate indirect evidence consisting of two sentences and three sentences, respectively [^7] We regulate the length of the generated evidence to control the influence of evidence length. The prompts used are detailed in Table 11 (index 7) in Appendix.. 4) Direct Evidence Combined with Indirect Evidence. We combine the direct evidence with the first sentence of the two-sentence indirect evidence to form a two-sentence evidence and with both sentences to form a three-sentence evidence.
@@
In Group 2, the MA Ratio ($R_{m}$) of direct evidence is slightly lower than that in Group 1. During the evidence generation, there are some questions for which ChatGPT can provide direct evidence but cannot produce indirect evidence. Removing these questions leads to a decrease in $R_{m}$ of direct evidence with one sentence, which implies that LLMs have a relatively high $R_{m}$ for the removed questions. But in general, the $R_{m}$ of direct evidence with one sentence in Group 1 is close to that in Group 2, so the results of Group 1 and Group 2 are still comparable.
Simple repetition of direct evidence is not always effective. Comparing direct evidence with one to three sentences, we observe similar $R_{m}$ and $R_{c}$ for LLaMA2-7B, LLaMA2-70B, and ChatGPT. For LLaMA3.2-3B, GPT-4, and Claude3.5, $R_{m}$ of direct evidence with two and three sentences decreases significantly. The results imply that the newer LLMs are receptive to evidence repeated multiple times.
Paraphrasing direct evidence is highly effective across all models and datasets. Comparing Direct Evidence with two and three sentences and Direct + Paraphrase with two and three sentences, respectively, we observe $R_{m}$ of the latter significantly decreases. For example, $R_{m}$ is reduced by more than half, comparing Direct + Paraphrase with three sentences with Direct Evidence with three sentences on the popQA dataset for all tested LLMs. The result implies that paraphrasing is an effective method to enhance the receptiveness of LLMs to external evidence.
Indirect Evidence improves LLMs’ receptiveness to CMAs, but less effectively than paraphrasing. Comparing Indirect Evidence with two and three sentences with Direct Evidence with one sentence, $R_{m}$ decreases for almost all LLMs, but the reduction is not significant compared to the Direct + Paraphrase evidence with two or three sentences. It implies that adding detailed information is less effective than paraphrasing direct evidence.
Combining Direct evidence with Indirect evidence generally enhances persuasiveness. Comparing Direct + Indirect evidence with Indirect Evidence, $R_{m}$ decreases except for LLaMA2-7B. For example, comparing Direct + Indirect with three sentences and Indirect Evidence with three sentences, $R_{m}$ has an obvious decrease. The result implies that adding direct evidence to indirect evidence is effective in improving LLMs’ receptiveness to CMAs.